OpenStudy (anonymous):

Inverse laplace trasform of s^2+2s+3/(s^2+2s+2) (s^2+2s+5)

OpenStudy (anonymous):

i'd say use partial fractions first

OpenStudy (anonymous):

when i factorise the bottom half i get complex number?

OpenStudy (anonymous):

i have attempted partial fractions .. still no look

OpenStudy (anonymous):

for the first part of the fraction i got (-1-i)(-1+i) , second part i got (-1-2i) (-1+2i)

OpenStudy (anonymous):

then i did a/(-1-i)+b(-1+i)+c/-1-2i+b/(-1+2i)

OpenStudy (anonymous):

see this\[\frac{s^2+2s+3}{(s^2+2s+2) (s^2+2s+5)}\]\[\frac{1}{2}\frac{2(s^2+2s+3)}{(s^2+2s+2) (s^2+2s+5)}\]\[\frac{1}{2}\frac{(s^2+2s+2)+(s^2+2s+5)-1}{(s^2+2s+2) (s^2+2s+5)}\]

OpenStudy (anonymous):

is that the corrrect way ?

OpenStudy (anonymous):

whats with the 1/2 and -1 where is that come from?

OpenStudy (anonymous):

yeah see\[\frac{1}{2}\frac{(s^2+2s+2)+(s^2+2s+5)-1}{(s^2+2s+2) (s^2+2s+5)}\]\[\frac{1}{2}({\frac{1}{s^2+2s+5}+\frac{1}{s^2+2s+2}}-\frac{1}{(s^2+2s+2) (s^2+2s+5)})\]

OpenStudy (anonymous):

right from this step you factorise ?